OpenStudy (anonymous):

Lim x^4-1/x^3-1 x->1

myininaya (myininaya):

Have you done any factoring?

OpenStudy (anonymous):

i tried but it just came back to the same spot.

myininaya (myininaya):

\[x^4-1=(x^2-1)(x^2+1)=(x-1)(x+1)(x^2+1)\] right?

myininaya (myininaya):

\[x^3-1=(x-1)(x^2+x+1)\]

OpenStudy (anonymous):

oooh yeah! thank you so much :))